# Large Language Models for Cryptocurrency Transaction Analysis: A Bitcoin Case Study

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
In addition, due to the input token limit of LLMs, how to efficiently feed larger graph data into LLMs to gain more information for potentially improving the quality of generated answers to various questions relevant to Bitcoin transaction graphs (e.g., address type prediction) continues to be an open question.
To bridge the gaps in applying LLMs to transaction graph analysis, we study Bitcoin networks and address three research questions:
@@
We adopt quantitative methods combined with qualitative analysis to answer those research questions. For RQ1, we investigate various graph representation formats and their feasibility for LLMs. To reduce the token consumption of raw graphs, we propose a novel representation format called LLM4TG based on the characteristics of LLMs. As for RQ2 and RQ3, we propose three levels for measuring the understandings of transaction graph:
@@
Contribution. Our work, for the first time, evaluate LLMs’ capabilities in analyzing transaction graphs for real-world cryptocurrencies, especially Bitcoin. We make the following contributions:
@@
    We apply our framework to evaluate LLMs’s capabilities (§4) under two datasets (BASD, BABD) and three mainstream models (GPT-3.5, GPT-4, GPT-4o).
@@ 3.1. Quick Overview
We present our framework in Fig. 1. The first step is to construct the Bitcoin transaction graph using the historical onchain data. The second step is to extract corresponding subgraphs (or the relevant subgraphs sampled by CETraS, §3.3) and graph features to Bitcoin addresses with labels. Finally, raw graphs (formatted as LLM4TG, §3.2) and graph features are input to LLMs with proper prompts according to the tasks on different understanding levels.
@@ 3.2. LLM4TG
Listing 1: Graph Model Definition
@@ 4.2. Level 1 - Foundational Metrics
Our results. Table 1 demonstrates that LLMs are excellent at node metrics. Accuracy for most metrics is between 98.50% and 100.00%. node_special_info_t is the exception. This may be due to the limited capability of LLMs to match many structurally similar data in transaction graphs. Compared with node metrics, however, for the global metrics, the accuracy significantly drops, ranging from 24.00% to 58.00%. Especially, we find that compared with the other metrics (35% to 58%) in global metrics, difference-related metrics are relatively low (24% to 34%). The reason for this may be that the capability of LLMs to calculate or compare is limited.
@@
Level 1 findings. For foundational metrics of transaction graphs, LLMs show high utility (98.50% to 100.00%) in node metrics and acceptable accuracy in several global metrics, especially for GPT-4o (44% to 58%). This illustrates that LLMs are excellent at obtaining specific node information in Bitcoin transaction graphs, but average at calculating or comparing data in the transaction graph.
@@ 4.3. Level 2 - Characteristic Overview
Our results. Based on our comprehensive empirical analysis of selected samples, the percentages of high-quality, average-quality, and low-quality cases for GPT-4 are 62.5%, 26.25% (inaccurate and invalid are 7.5% and 18.75%), and 11.25%; for GPT-4o are 82.5%, 13.75% (inaccurate and invalid are 12.5% and 1.25%), and 3.75%. That means the proportions of meaningful responses for GPT-4 and GPT-4o are 70% and 95%. We select two examples, i.e.,bc1qah111bc1qahe54yxl33clnwdtleuh4cw0fw4df62t0tnuk2 and bc1q4w222bc1q4w090gzj7m9hy39l8e48gzz4uvdfn2x9y9g70h, to ensure each category is included to further illustrate our results, combined with visualization by Gephi. We apply the ID of address node $n_{0}$ to represent the corresponding subgraph.
@@
Level 2 findings. Despite some inaccuracies responses exist, LLMs extract numerous beneficial features for transaction analysis, achieving 70% and 95% for GPT-4 and GPT4o, respectively. This indicates that LLMs have an exceptional characteristic overview understanding capability, especially GPT-4o, highlighting their utility in characteristics identification within Bitcoin transaction graphs.
@@ 4.4. Level 3 - Contextual Interpretation
Graph feature-based classification. For the graph feature-based classification tasks, we employ five labeled subgraphs per category, represented by subgraph IDs and their corresponding features, along with 500 randomly chosen unlabeled subgraphs with features. Except for inputting the values of these features to LLMs, we also input the corresponding descriptions. We select the ten most important features in BABD (10375557, ) (Table 2). Our analysis focuses on the accuracy, macro precision, recall, and F1 scores of LLMs and their performance in each class.
The overall accuracy of LLMs (Fig.3(a)) is between 39.83% to 46.07%, which is not impressive, as well as the macro precision, macro recall, and macro F1 score. However, this is acceptable since the reference samples are significantly few. In contrast, the top-3 accuracy achieves high, ranging from 67.20% to 71.02%. This suggests that while achieving exact classification might be challenging due to the limited data, LLMs can identify the correct classes within their top three predictions.
Beyond overall metrics, class-specific results are displayed in Fig.4. GPT-4 and GPT-4o show strong performance in mining pool, achieving 80% and 95% precisions; They also reach extremely high recalls for darknet market which are 98.46% and 96.92%. For the majority of classes, GPT-4 and GPT-4o show better performance, though in some classes GPT-3 has similar or higher performance (e.g., recall for money laundering and precision for darknet market).
@@
For various LLMs, GPT-4o showcases slightly higher overall accuracy, macro precision, macro recall, and macro F1 score than GPT-3.5 and GPT-4, which is generally below 10%. Thus, the update and optimization of LLMs may not achieve huge enhancement of graph feature-based classification tasks. Furthermore, for each class, GPT-4o has fewer categories with precision, recall, or F1 scores below 10% compared with GPT-3.5 and GPT-4. To sum up, GPT-4o demonstrates mildly superior performance over GPT-3.5 and GPT-4 in graph feature-based classification tasks, however, the improvements are generally minimal.
@@
For the specific class as shown in Figure 5, GPT-4o performs better than GPT-4 in almost all metrics, in some classes such as darknet market, the differences of F1 scores are even close to 80%. One remarkable exception is that both GPT-4 and GPT-4o recorded zero scores in all metrics on blackmail; However, for GPT-4, classes including gambling, and Ponzi scheme recorded zero scores across all metrics, reflecting a complete failure in identifying relevant instances. This illustrates the optimization of LLMs may be effective for raw graph-based classification tasks.
@@ 5. Discussion
    Answer to RQ2. We propsed a three-level framework for better understanding the graphs. For each level, we have established clear standards and conducted experiments to validate their practical application. The results underscore the framework’s effectiveness in enhancing the analytical capabilities of LLMs in this domain.
@@
    Data representation and features. The type of data used (e.g., graph features versus raw graphs) and the availability of labeled samples affect performance. Graph feature-based data is more efficient in terms of LLM token consumption and can improve classification effectiveness with additional features and labels.
6. Related Work
Empirical analysis in cryptocurrency. Empirical analysis plays an important role in understanding the dynamics of the cryptocurrency ecosystem (wang2024cryptocurrency, ; huang2023miracle, ; wang2022empirical, ) and the behaviors of addresses/entities (chen2020understanding, ; gao2020tracking, ). For instance, Tovanich et al. (tovanich2021empirical, ) and Hou et al. (hou2020study, ) reveal that factors such as payout schemes and pool fees influence miners’ behaviors in Bitcoin mining pools, and then impact the overall system performance.
Empirical analysis can also aid in identifying cryptocurrency scams. Li et al. (li2023double, ) used CryptoScamTracker to analyze cryptocurrency giveaway scams. He et al. (he2023txphishscope, ) developed TxPhish to detect Ethereum scams in which users lured by high profits to fake websites are tricked into signing transactions that allow scammers to steal their crypto assets. Gomez et al. (gomez2022watch, ) explored bidirectional transactions to map cybercrime networks. Wu et al. (10371347, ) focused on identifying Ethereum-based money laundering via asset flows. Chen et al. (chen2022cryptocurrency, ) combined on-chain and off-chain data to detect fake trading volumes of famous exchanges Huobi and Binance.
  • $\bullet$
Nerurkar et al. (nerurkar2021dissecting, ) and Serena et al. (serena2022cryptocurrencies, ) drew parallels between cryptocurrency systems and other complex systems and identified characteristics such as small-world property, indicating most nodes in the graph are not neighbors but most of them can be reached by every other within a few hops. Moreover, Tao et al. (tao2021complex, ) employed an innovative random walk with a flying-back sampling method on Bitcoin transaction graphs, uncovering phenomena such as the non-rich-club effect, i.e., that high-degree nodes are not more interconnected among themselves than with lower-degree nodes. Guo et al. (guo2019graph, ) analyzed Ethereum transaction graphs, revealing heavy-tailed property in transaction networks, i.e., the majority of nodes have a relatively low degree while a small number of nodes have a very high degree. To mitigate the issue of money laundering on blockchain networks, DenseFlow framework, proposed by Lin et al. (lin2024denseflow, ), uses dense subgraphs and the maximum flow algorithm to trace laundering activities. This approach improves precision compared to existing methods on Ethereum, demonstrating the effectiveness of network analysis in combating money laundering.
  • $\bullet$
Chaudhari et al. (chaudhari2021towards, ) studied utilizing temporal features to detect Bitcoin address behavioral changes and identify money laundering activities. The proposed approaches based on the decision tree (DT) by Rathore et al. (rathore2022mixers, ) show high accuracy rates in detecting illicit activities and phishing scams in cryptocurrencies. Wahrstatter et al. (wahrstatter2023improving, ) also contributed by enhancing the detection of criminal activities in Bitcoin transactions using unsupervised learning. Additionally, various machine learning methods, such as random forest (RF), multilayer perceptron (MLP), and graph neural network (GNN), are applied to benchmark cryptocurrency datasets, including the Elliptic Data Set by Weber et al. (weber2019anti, ) and datasets by Xiang et al. (10375557, ; xiang2022leveraging, ). Besides, Gai et al. (gai2023blockchain, ) proposed a transformer-based anomaly detection model BlockGPT for the Ethereum network that demonstrated acceptable utility.
LLMs in graph analysis. Several studies introduced LLMs to graph analysis, using LLM as a classifier and GNN enhancement (li2023survey, ). Wang et al. (wang2023can, ; yu2023predicting, ) evaluated the basic capabilities of LLMs in natural language graph problem-solving. Both studies showed limitations of LLMs, particularly in solving complex graph structures and tasks. Complementing these insights, Liu and Wu (liu2023evaluating, ) and Hu et al. (hu2023beyond, ) assessed the performance of LLMs in graph data analysis and prediction, compared them with specialized GNNs. Likewise, Sui et al. (sui2023gpt4table, ) and Jiang et al. (jiang2023structgpt, ) explored the effectiveness of LLMs including GPT-3.5 and GPT-4 in processing structured data, such as tables and various structured data types, introducing innovative prompting methods for performance enhancement.
Besides, Das et al. (das2023modality, ), Chen et al. (chen2023exploring, ; chen2023labelf, ), and Guo et al. (guo2023gpt4graph, ) adopted a different approach by integrating LLMs with graph data, focusing on graph structure analysis, node classification, and a range of graph processing tasks. These studies investigated the potential and limitations of LLMs in more specialized and advanced graph analysis applications, offering new insights and directions for future research in LLMs and graph data analysis. Moreover, Sun et al. (sun2023head, ) uniquely studied the factual knowledge of LLMs, providing a broader perspective on their comprehension capabilities, especially for lesser-known entities and facts.
  • $\bullet$
    Existing research mainly focuses on knowledge graphs and randomly generated graphs (das2023modality, ; hu2023beyond, ; liu2023evaluating, ; wang2023can, ). However, how to measure the LLMs’ ability to understand and analyze real-world cryptocurrency transaction graphs is still unresolved.
  • $\bullet$
    Common graph representation formats, such as GEXF and GraphML, are not ideally suited for LLMs due to their inherent space constraints. This limitation explains why recent studies have focused exclusively on testing LLMs with smaller graphs (das2023modality, ; wang2023can, ) (e.g., graphs containing ten nodes or marginally more (wang2023can, )).
  • $\bullet$
@@ 7. Conclusion
This work evaluates the capabilities of LLMs in analyzing Bitcoin transaction graphs. We introduced a three-level framework along with two key innovations: the LLM4TG format to enhance readability, and the CETraS algorithm to reduce graph sizes. Our experiments demonstrate satisfactory accuracy in foundational metrics and effectiveness in obtaining useful overview characteristics.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Intrusion/anomaly detection and malware mitigationccs: Applied computing Network forensicsccs: Computing methodologies Machine learning
@@ 1. Introduction
In addition, due to the input token limit of LLMs, how to efficiently feed larger graph data into LLMs to gain more information for potentially improving the quality of generated answers to various questions relevant to Bitcoin transaction graphs (e.g., address type prediction) continues to be an open question. To bridge the gaps in applying LLMs to transaction graph analysis, we study Bitcoin networks and address three research questions:
@@
We adopt quantitative methods combined with qualitative analysis to answer those research questions. For RQ1, we investigate various graph representation formats and their feasibility for LLMs. To reduce the token consumption of raw graphs, we propose a novel representation format called LLM4TG based on the characteristics of LLMs. As for RQ2 and RQ3, we propose three levels for measuring the understanding of transaction graph:
@@
Contribution. Our work, for the first time, evaluates LLMs’ capabilities in analyzing transaction graphs for real-world cryptocurrencies, especially Bitcoin. We make the following contributions:
@@
    We applied our framework to evaluate LLMs’ capabilities in real-world cryptocurrency transaction analysis using two datasets (BASD, BABD) extracted from the Bitcoin ledger and three mainstream models (GPT-3.5, GPT-4, GPT-4o), providing both quantitative assessment and qualitative analysis of their performance.
@@ 3.1. Quick Overview
We present our framework in Fig. 1. The first step is to construct the Bitcoin transaction graph using the historical on-chain data. The second step is to extract corresponding subgraphs (or the relevant subgraphs sampled by CETraS, §3.3) and graph features to Bitcoin addresses with labels. Finally, raw graphs (formatted as LLM4TG, §3.2) and graph features are input to LLMs with proper prompts according to the tasks on different understanding levels.
@@ 3.2. LLM4TG
Listing 1: Graph Representation Definition
@@ 4.2. Level 1 - Foundational Metrics
Our results. Table 1 demonstrates that LLMs are excellent at node metrics. Accuracy for most metrics is between 98.50% and 100.00%. node_special_info_t is the exception. This may be due to the limited capability of LLMs to match many structurally similar data in transaction graphs. Compared with node metrics, however, for the global metrics, the accuracy significantly drops, ranging from 24.00% to 58.00%. Especially, we find that compared with the other metrics (35.00% to 58.00%) in global metrics, difference-related metrics are relatively low (24.00% to 34.00%). The reason for this may be that the capability of LLMs to calculate or compare is limited.
@@
Level 1 findings. For foundational metrics of transaction graphs, LLMs show high utility (98.50% to 100.00%) in node metrics and acceptable accuracy in several global metrics, especially for GPT-4o (44.00% to 58.00%). This illustrates that LLMs are excellent at obtaining specific node information in Bitcoin transaction graphs, but average at calculating or comparing data in the transaction graph.
@@ 4.3. Level 2 - Characteristic Overview
Our results. Based on our comprehensive empirical analysis of selected samples, the percentages of high-quality, average-quality, and low-quality cases for GPT-4 are 62.50%, 26.25% (inaccurate and invalid are 7.50% and 18.75%), and 11.25%; for GPT-4o are 82.50%, 13.75% (inaccurate and invalid are 12.50% and 1.25%), and 3.75%. That means the proportions of meaningful responses for GPT-4 and GPT-4o are 70.00% and 95.00%. We select two examples, i.e.,bc1qah111bc1qahe54yxl33clnwdtleuh4cw0fw4df62t0tnuk2 and bc1q4w222bc1q4w090gzj7m9hy39l8e48gzz4uvdfn2x9y9g70h, to ensure each category is included to further illustrate our results, combined with visualization by Gephi. We apply the ID of address node $n_{0}$ to represent the corresponding subgraph.
@@
Level 2 findings. Despite some inaccuracies responses exist, LLMs extract numerous beneficial features for transaction analysis, achieving 70.00% and 95.00% for GPT-4 and GPT4o, respectively. This indicates that LLMs have an exceptional characteristic overview understanding capability, especially GPT-4o, highlighting their utility in characteristics identification within Bitcoin transaction graphs.
@@ 4.4. Level 3 - Contextual Interpretation
Graph feature-based classification. For the graph feature-based classification tasks, we employ five labeled subgraphs per category, where each subgraph is represented by its starting Bitcoin address (i.e., $n_{0}$) and its corresponding features, along with 500 randomly chosen unlabeled subgraphs with features. Except for inputting the values of these features to LLMs, we also input the corresponding descriptions. We select the ten most important features in BABD (10375557, ) (Table 2). Our analysis focuses on the accuracy, macro precision, recall, and F1 scores of LLMs and their performance in each class.
The overall accuracy of LLMs (Fig.3(a)) is between 39.83% to 46.07%, which is not impressive, as well as the macro precision, macro recall, and macro F1 score. However, this is acceptable since the reference samples are significantly few. In contrast, the top-3 accuracy achieves high, ranging from 67.20% to 71.02%. This suggests that while achieving exact classification might be challenging due to the limited data, LLMs can achieve relatively high accuracy in identifying the correct classes within their top three predictions.
Beyond overall metrics, class-specific results are displayed in Fig.4. GPT-4 and GPT-4o show strong performance in mining pool, achieving 80.00% and 95.00% precisions; They also reach extremely high recalls for darknet market which are 98.46% and 96.92%. For the majority of classes, GPT-4 and GPT-4o show better performance, though in some classes GPT-3 has similar or higher performance (e.g., recall for money laundering and precision for darknet market).
@@
Overall, GPT-4o achieves the highest top-3 accuracy (71.02%) and recall (47.45%), while also outperforming GPT-3.5 and GPT-4 in macro precision. However, its accuracy remains lower than GPT-4, and its F1 score, while surpassing GPT-3.5, is slightly lower than GPT-4. Nevertheless, GPT-4o demonstrates better stability, with fewer categories having macro precision, recall, or F1 scores below 10%. These results suggest that despite updates leading to marginal improvements in graph feature-based classification, GPT-4o still offers mild but limited advantages over GPT-3.5 and GPT-4.
@@
For the specific class as shown in Figure 5, GPT-4o performs better than GPT-4 in almost all metrics, in some classes such as darknet market, the differences of F1 scores are even close to 80%. One remarkable exception is that both GPT-4 and GPT-4o recorded zero scores in all metrics on blackmail; However, for GPT-4, classes including gambling, and Ponzi scheme recorded zero scores across all metrics, reflecting a huge failure in identifying relevant instances.
To sum up, GPT-4o achieves significantly higher performance than GPT-4 across all metrics in raw graph-based tasks and outperforms it in most specific classes. This demonstrates that optimizing LLMs may effectively enhance raw graph-based classification tasks.
@@ 5. Discussion
    Answer to RQ2. We proposed a three-level framework for better understanding transaction graphs. For each level, we have established clear standards and conducted experiments to validate their practical application. The results underscore the framework’s effectiveness in enhancing the analytical capabilities of LLMs in this domain.
@@
    Data representation and features. The type of data used (e.g., graph features versus raw graphs) and the availability of labeled samples affect performance. Graph feature-based data is more efficient in terms of LLM token consumption and may improve classification effectiveness with additional features and labels.
@@ 6. Conclusion
This work evaluates LLMs’ capabilities in analyzing Bitcoin transaction graphs. We introduced a three-level framework along with two key innovations: LLM4TG format to enhance readability and reduce graph sizes, and CETraS algorithm to optimize graph simplification. Our experiments demonstrate satisfactory accuracy in foundational metrics, effectiveness in obtaining useful overview characteristics, and solid top-3 accuracy in classification tasks. These findings highlight the significant potential of LLMs and establish a foundation for their broader application in cryptocurrency analysis.
Acknowledgment
This research is supported by OpenAI Researcher Access Program.
@@ Appendix A LLMs Responses
Appendix B Related Work
Empirical analysis in cryptocurrency. Empirical analysis plays an important role in understanding the dynamics of the cryptocurrency ecosystem [27, 28, 29] and the behaviors of addresses/entities [14, 30]. For instance, Tovanich et al. [31] and Hou et al. [32] reveal that factors such as payout schemes and pool fees influence miners’ behaviors in Bitcoin mining pools, and then impact the overall system performance.
Empirical analysis can also aid in identifying cryptocurrency scams. Li et al. [33] used CryptoScamTracker to analyze cryptocurrency giveaway scams. He et al. [34] developed TxPhish to detect Ethereum scams in which users lured by high profits to fake websites are tricked into signing transactions that allow scammers to steal their crypto assets. Gomez et al. [35] explored bidirectional transactions to map cybercrime networks. Wu et al. [36] focused on identifying Ethereum-based money laundering via asset flows. Chen et al. [37] combined on-chain and off-chain data to detect fake trading volumes of famous exchanges Huobi and Binance.
  • $\bullet$
Nerurkar et al. [38] and Serena et al. [39] drew parallels between cryptocurrency systems and other complex systems and identified characteristics such as small-world property, indicating most nodes in the graph are not neighbors but most of them can be reached by every other within a few hops. Moreover, Tao et al. [40] employed an innovative random walk with a flying-back sampling method on Bitcoin transaction graphs, uncovering phenomena such as the non-rich-club effect, i.e., that high-degree nodes are not more interconnected among themselves than with lower-degree nodes. Guo et al. [16] analyzed Ethereum transaction graphs, revealing heavy-tailed property in transaction networks, i.e., the majority of nodes have a relatively low degree while a small number of nodes have a very high degree. To mitigate the issue of money laundering on blockchain networks, DenseFlow framework, proposed by Lin et al. [41], uses dense subgraphs and the maximum flow algorithm to trace laundering activities. This approach improves precision compared to existing methods on Ethereum, demonstrating the effectiveness of network analysis in combating money laundering.
  • $\bullet$
Chaudhari et al. [42] studied utilizing temporal features to detect Bitcoin address behavioral changes and identify money laundering activities. The proposed approaches based on the decision tree (DT) by Rathore et al. [43] show high accuracy rates in detecting illicit activities and phishing scams in cryptocurrencies. Wahrstatter et al. [44] also contributed by enhancing the detection of criminal activities in Bitcoin transactions using unsupervised learning. Additionally, various machine learning methods, such as random forest (RF), multilayer perceptron (MLP), and graph neural network (GNN), are applied to benchmark cryptocurrency datasets, including the Elliptic Data Set by Weber et al. [15] and datasets by Xiang et al. [13, 25]. Besides, Gai et al. [45] proposed a transformer-based anomaly detection model BlockGPT for the Ethereum network that demonstrated acceptable utility.
LLMs in graph analysis. Several studies introduced LLMs to graph analysis, using LLM as a classifier and GNN enhancement [20]. Wang et al. [11, 17] evaluated the basic capabilities of LLMs in natural language graph problem-solving. Both studies showed limitations of LLMs, particularly in solving complex graph structures and tasks. Complementing these insights, Liu and Wu [46] and Hu et al. [10] assessed the performance of LLMs in graph data analysis and prediction, compared them with specialized GNNs. Likewise, Sui et al. [47] and Jiang et al. [48] explored the effectiveness of LLMs including GPT-3.5 and GPT-4 in processing structured data, such as tables and various structured data types, introducing innovative prompting methods for performance enhancement.
Besides, Das et al. [49], Chen et al. [50, 51], and Guo et al. [9] adopted a different approach by integrating LLMs with graph data, focusing on graph structure analysis, node classification, and a range of graph processing tasks. These studies investigated the potential and limitations of LLMs in more specialized and advanced graph analysis applications, offering new insights and directions for future research in LLMs and graph data analysis. Moreover, Sun et al. [52] uniquely studied the factual knowledge of LLMs, providing a broader perspective on their comprehension capabilities, especially for lesser-known entities and facts.
  • $\bullet$
    Existing research mainly focuses on knowledge graphs and randomly generated graphs [49, 10, 46, 11]. However, how to measure the LLMs’ ability to understand and analyze real-world cryptocurrency transaction graphs is still unresolved.
  • $\bullet$
    Common graph representation formats, such as GEXF and GraphML, are not ideally suited for LLMs due to their inherent space constraints. This limitation explains why recent studies have focused exclusively on testing LLMs with smaller graphs [49, 11] (e.g., graphs containing ten nodes or marginally more [11]).
  • $\bullet$